### v1

# StochasticBarrier.jl: A Toolbox for Stochastic Barrier Function Synthesis

Abstract.
We present StochasticBarrier.jl, an open-source Julia-based toolbox for generating Stochastic Barrier Functions (SBFs) for safety verification of discrete-time stochastic systems with additive Gaussian noise. StochasticBarrier.jl certifies linear, polynomial, and piecewise affine (PWA) systems. The latter enables verification for a wide range of system dynamics, including general nonlinear types. The toolbox implements a Sum-of-Squares (SOS) optimization approach, as well as methods based on piecewise constant (PWC) functions. For SOS-based SBFs, StochasticBarrier.jl leverages semi-definite programming solvers, while for PWC SBFs, it offers three engines: two using linear programming (LP) and one based on gradient descent (GD). Benchmarking StochasticBarrier.jl against the state-of-the-art shows that the tool outperforms existing tools in computation time, safety probability bounds, and scalability across over 30 case studies. Compared to its closest competitor, StochasticBarrier.jl is up to four orders of magnitude faster, achieves significant safety probability improvements, and supports higher-dimensional systems.
Barrier Certificates, Stochastic Systems, Probabilistic Safety, Convex Optimization, Formal Verification
1. Introduction
Stochastic barrier functions (SBFs) are a class of tools used for the formal verification of stochastic systems (Prajna et al., 2007; Santoyo et al., 2021). These functions establish a lower bound on the probability of a system remaining within a safe set, using martingale theory (Santoyo et al., 2021; Mazouz et al., 2022). A key advantage of SBFs over other verification methods is their capacity to analyze the time evolution of a system through a set of static constraints, eliminating the need to propagate system uncertainty. Instead, SBF synthesis methods solve a functional optimization problem, for which various approaches have been developed based on, e.g., Sum-of-Squares (SOS) polynomials (Prajna et al., 2007; Santoyo et al., 2021; Mazouz et al., 2022) and piecewise (PW) functions (Mazouz et al., 2024a, b; Mathiesen et al., 2024b). Despite their potential, however, there has been limited effort to develop robust and efficient tools for SBF synthesis, restricting their broader application in practical fields such as robotics and autonomous vehicles.
In this paper, we address this gap by introducing an open-source Julia-based toolbox that generates SBFs for discrete-time stochastic systems with additive Gaussian noise - StochasticBarrier.jl   [^1] https://github.com/aria-systems-group/StochasticBarrier.jl.. The toolbox supports a wide range of dynamics, namely, linear, polynomial, and PW affine (PWA) inclusions, the latter enabling the verification of non-linear, non-polynomial systems. The established SOS optimization method is implemented, as well as more recent approaches based on PW Constant (PWC) functions. For SOS optimization, existing semi-definite programming (SDP) solvers are leveraged. For PWC-based SBFs, three engines are included in StochasticBarrier.jl: two that utilize Linear Programming (LP) and one based on Gradient Descent (GD), as introduced in (Mazouz et al., 2024a).
The toolbox is developed in Julia, a modern programming language designed for the scientific community. Julia combines the interactivity and dynamic typing of scripting languages, facilitating fast prototyping, with the high performance of compiled languages (Perkel and others, 2019). Also, Julia’s support for parametric typing allows for customizable precision in numerical computations (Mathiesen et al., 2024a). We evaluate StochasticBarrier.jl on various benchmarks and compare its performance against two state-of-the-art tools: PRoTECT (Wooding et al., 2024) and StochasticBarrierFunctions (Santoyo et al., 2021), which, to the best of our knowledge, are the only available SBF tools for discrete-time stochastic systems. These existing tools exclusively support SOS-based approaches. In addition, the three PWC-SBF methods are compared against the SOS-method. Benchmarks include over 30 case studies across eight different systems with linear, polynomial, and PWA inclusion dynamics, ranging from one to six dimensions. The results demonstrate that StochasticBarrier.jl outperforms existing tools in three metrics: computation time, safety probability bounds, and scalability. Compared to its closest competitor, StochasticBarrier.jl is up to $1000\times$ faster for successful synthesis, improves safety probability bounds from 0 to $1.0$ in some cases, and handles systems with twice the dimensionality.
2. Theoretical Framework
We provide an overview of the SBF theoretical framework.
2.1. Problem
Consider a stochastic process defined by difference equation
(1) $\mathbf{x}_{k+1}=f(\mathbf{x}_{k})+\mathbf{w}_{k},$
where $\mathbf{x}_{k}\in\mathbb{R}^{n}$ is the state, $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is the vector field representing the one-step dynamics of System (1), and $\mathbf{w}_{k}\in\mathbb{R}^{n}$ is an independent and identically distributed random variable, representing noise. The probability distribution of $\mathbf{w}_{k}$ is assumed to be zero-mean Gaussian, i.e., $\mathbf{w}_{k}\sim p_{\mathbf{w}}=\mathcal{N}(0,\Sigma)$, where $\Sigma=\mathrm{diag}\!\left(\sigma_{1}^{2},\ldots,\sigma_{n}^{2}\right)$. We define the stochastic kernel $T(X\mid x):=\int_{\mathbb{R}^{\mathrm{w}}}\mathbf{1}_{X}(f(x)+w)p_{\mathbf{w}}(dw)$. It follows from this definition of $T$ that, given an initial condition $\mathbf{x}_{0}=x_{0}\in\mathbb{R}^{n}$, $\mathbf{x}_{k}$ is a Markov process (Bertsekas and Shreve, 2004). Let $X_{\mathrm{s}}\subset\mathbb{R}^{n}$ be a bounded set representing the safe set, $X_{0}\subseteq X_{\mathrm{s}}$ be the initial set, and $N\in\mathbb{N}$ be the time horizon. Then, probabilistic safety is defined as $P_{\mathrm{s}}(X_{\mathrm{s}},X_{0},N)=\inf_{x_{0}\in X_{0}}\mathrm{Pr}^{x_{0}}[\forall k\in\{0,...,N\},\mathbf{x}_{k}\in X_{\mathrm{s}}].$
To reason about $P_{\mathrm{s}}(X_{\mathrm{s}},X_{0},N)$, Stochastic Barrier Function (SBFs) can be used (Santoyo et al., 2021; Mazouz et al., 2022). Function $B:\mathbb{R}^{n}\to\mathbb{R}_{\geq 0}$ is a SBF for System (1) if there exist scalars $\eta,\beta\geq 0$, such that
(2a) $B(x)\geq 1\qquad$ $\forall x\in\mathbb{R}^{n}\setminus X_{s},$
(2b) $B(x)\leq\eta\qquad$ $\forall x\in X_{0},$
(2c) $\mathbb{E}[B(f(x)+\mathbf{w}))]\leq B(x)+\beta\qquad$ $\forall x\in X_{s},$
which guarantees a lower-bound for the probabilistic safety
(3) $P_{\mathrm{s}}(X_{\mathrm{s}},X_{0},N)\geq 1-(\eta+N\beta).$
The goal of the SBF synthesis problem is to find a $B(x)$ that maximizes $P_{s}$, subject to Conditions (2a)-(2c). StochasticBarrier.jl solves this problem by computing a $B$ along with a $P_{s}$ lower bound.
From a computational perspective, solving this functional optimization problem is non-trivial. Several methods have been developed to ensure the convexity of the problem. One well-established approach assumes $B$ is a polynomial of a given degree, which allows the problem to be formulated as a SOS optimization, ultimately leading to a semi-definite program (Santoyo et al., 2021; Prajna, 2006; Mazouz et al., 2022). A more recent approach, introduced in (Mazouz et al., 2024a), assumes $B$ is a piecewise constant (PWC) function, reducing the optimization problem to a linear program. Both methods are implemented in StochasticBarrier.jl.
2.2. Implemented SBF Methods
StochasticBarrier.jl supports polynomial (SOS) and PWC class-es of SBFs for System (1), where $f$ is affine, polynomial, or bounded by piecewise affine (PWA) functions.
2.2.1. Sum-of-Squares (SOS)
A multivariate polynomial $\lambda(x)$ with $x\in\mathbb{R}^{n}$ is an SOS polynomial if there exist polynomials $\lambda_{i}$, $i=1,\ldots,r$, for $r\in\mathbb{N}$}, such that $\lambda(x)=\sum_{i=1}^{r}\lambda_{i}^{2}(x).$ If $\lambda(x)$ is SOS, then $\lambda(x)\geq 0$ for all $x\in\mathbb{R}^{n}$. The set of SOS polynomials is denoted by $\Lambda(x)$, and a vector of SOS polynomials by $\mathcal{L}(x)$. Under the assumption that $f$ in System (1) is a polynomial of degree one or higher, and let $X_{0},X_{s}$ be compact semi-algebraic sets, then, the following theorem (Santoyo et al., 2021) is implemented in StochasticBarrier.jl to reduce the SOS-SBF synthesis problem to a convex program.
Theorem 2.1 (SOS-SBF for Polynomial Systems (Santoyo et al., 2021)).
Let the safe set $X_{\mathrm{s}}=\{x\in\mathbb{R}^{n}\mid h_{\mathrm{s}}(x)\geq 0\}$, initial set $X_{\mathrm{0}}=\{x\in\mathbb{R}^{n}\mid h_{\mathrm{0}}(x)\geq 0\}$, and unsafe set $X_{\mathrm{u}}=\mathbb{R}^{n}\setminus X_{\mathrm{s}}=\{x\in\mathbb{R}^{n}\mid h_{\mathrm{u}}(x)\geq 0\}$, where $h_{s},h_{0},h_{u}$ are vectors of polynomials, be given. Then, a SBF $B(x)\in\Lambda(x)$ for System (1) is obtained by solving the following SOS optimization problem
(4a) $\min_{\beta,\eta\geq 0}\quad\eta+N\beta\qquad\text{subject to:}$
(4b) $\hskip 11.38109ptB(x)\in\Lambda(x),$
(4c) $\hskip 1.42262pt-B(x)-\mathcal{L}^{T}_{\mathrm{0}}(x)h_{\mathrm{0}}(x)+\eta\in\Lambda(x),$
(4d) $\hskip 11.38109ptB(x)-\mathcal{L}^{T}_{\mathrm{u}}(x)h_{\mathrm{u}}(x)-1\in\Lambda(x),$
(4e) $\hskip 1.42262pt-\mathbb{E}[B(f(x)+\mathbf{w})]+B(x)+\beta-\mathcal{L}^{T}_{s}(x)h_{s}(x)\in\Lambda(x),$
where $\mathcal{L}_{0}$, $\mathcal{L}_{h}$, and $\mathcal{L}_{s}$ are vectors of SOS polynomials of appropriate dimension. A valid solution to this optimization guarantees safety probability in Eq. (3).
PWA Inclusion Dynamics
Additionally to polynomial $f$, non-polynomial $f$ are also supported in StochasticBarrier.jl by supplying piecewise affine (PWA) bounds of $f$. In particular, given a partition $Q=\{q_{1},\ldots,q_{|Q|}\}$ of the safe set $X_{\mathrm{s}}$ for every $q\in Q$, StochasticBarrier.jl supports $f$ expressed via bounds
(5) $\underline{f}_{q}(x)=\underline{A}_{q}x+\underline{b}_{q}\quad\text{and}\quad\overline{f}_{q}(x)=\overline{A}_{q}x+\overline{b}_{q}$
such that $\underline{f}_{q}(x)\leq f(x)\leq\overline{f}_{q}(x)$ for every $x\in q$. The following theorem shows a convex optimization framework for synthesis of SOS-SBF for such dynamics.
Theorem 2.2 (SOS-SBF for PWA Inclusion Systems (Mazouz et al., 2022)).
Consider function $B(x)$ and sets $X_{0}$, $X_{s}$, $X_{u}$ as in Theorem 2.1 and $X_{s}$-partition $Q=\{q_{1},\ldots,q_{|Q|}\}$, where each $q\in Q$ is a semi-algebraic set $q=\{x\in\mathbb{R}^{n}\mid h_{{q}}(x)\geq 0\}$. Given PWA functions $\underline{f}_{q}$ and $\overline{f}_{q}$ over $Q$ as defined in Eq. (5), an SOS SBF $B(x)$ for System (1) with general $f(x)\in[\underline{f}_{q}(x),\overline{f}_{q}(x)]$ can be obtained by solving the SOS optimization problem in (4) and replacing Constraint (4e) with $\mathbb{E}[B(y+{w})\mid x]+B(x)+\beta-\mathcal{L}^{T}_{q,x}(x)h_{q}(x)-\mathcal{L}^{T}_{q,y}(x)\big((\overline{f}_{q}(x)-y)\odot(y-\underline{f}_{q}(x))\big)\in\Lambda(x,y),$ for all $q\in Q$, where $\odot$ is the Schur product.
In StochasticBarrier.jl, the implementation of partition $Q$ is based on a grid, where each $q\in Q$ is defined as a hyper-rectangle. Then, provided partition $Q$ and PWA functions $\underline{f}_{q}$ and $\overline{f}_{q}$, the tool performs the above optimization.
2.2.2. PWC Method
StochasticBarrier.jl  also supports the class of PWC-SBFs. Consider a partition $Q=\{q_{1},\ldots,q_{|Q|}\}$ of $X_{s}$ such that vector field $f(x)$ is continuous in each region $q\in Q$. Define PWC function $B:\mathbb{R}^{n}\to\mathbb{R}$ as
(6) $B(x)=b_{i}\;\text{ if }\;x\in q_{i},\text{ otherwise }B(x)=1.$
where $b_{i}\in\mathbb{R}$. Further, let $q_{u}=X_{u}$ and denote the bounds on the transition kernel $T$ from partition $q_{i}\in Q$ by, for all $j\in\{1,\ldots,|Q|,u\}$,
(7) $\underline{p}_{ij}=\inf_{x\in q_{i}}T(q_{j}\mid x),\quad\text{and}\quad\overline{p}_{ij}=\sup_{x\in q_{i}}T(q_{j}\mid x).$
Then, the set of feasible values for $T$ from $q_{i}$ is given by $\mathcal{P}_{i}=\big\{p_{i}=(p_{i1},\ldots,p_{i|Q|},p_{i\mathrm{u}})\in[0,1]^{|Q|+1}$   s.t.   $\sum_{j=1}^{|Q|}p_{ij}+p_{i\mathrm{u}}=1,\;\underline{p}_{ij}\leq p_{ij}\leq\overline{p}_{ij}\;\;\;\forall j\in\{1,\ldots,|Q|,\mathrm{u}\}\big\}.$ The following theorem formalizes the optimization problem for PWC class of SBFs for System (1).
Theorem 2.3 (PWC-SBF (Mazouz et al., 2024a)).
Consider System (1) with safe set $X_{s}$, initial set $X_{0}$, and partition $Q=\{q_{1},\ldots,q_{|Q|}\}$ of $X_{s}$. Let $\mathcal{B}_{Q}$ be the class of PWC functions in the form of Eq. (6) and $\mathcal{P}=\mathcal{P}_{1}\times\ldots\times\mathcal{P}_{|Q|}$, where each $\mathcal{P}_{i}$ is the set of feasible transitions. Then, $B^{*}\in\mathcal{B}_{Q}$ is a PWC-SBF if $B^{*}$ is a solution to
$B^{*}=\arg\min_{B\in\mathcal{B}_{Q}}\;\max_{(p_{i})_{i=1}^{|Q|}\in\mathcal{P}}\;\eta+N\beta$     subject to,
$\forall i\in\{1,\ldots,Q\},\;\;b_{i}\geq 0,\;\;\sum_{j=1}^{Q}b_{j}\cdot p_{ij}+p_{i\mathrm{u}}\leq b_{i}+\beta_{i},\;\;0\leq\beta_{i}\leq\beta,$
and for all $i:X_{i}\cap X_{0}\neq\emptyset$,    $b_{i}\leq\eta$. Then, $B^{*}$ guarantees the safety probability in Eq. (3).
Work (Mazouz et al., 2024a) introduces three methods for synthesizing PWC SBFs: (i) an exact duality-based approach formulated as a zero-duality-gap Linear Program (LP), (ii) an exact Counter-Example Guided Inductive Synthesis (CEGIS) method that alternates between generating candidate barriers and counter-examples via two LPs, and (iii) a gradient descent (GD) method offering improved scalability and memory efficiency. StochasticBarrier.jl implements all three methods for System (1) with linear and PWA inclusion dynamics.
3. Overview of StochasticBarrier.jl
Table 1. Overview of SBF synthesis methods implemented in StochasticBarrier.jl for different classes of dynamics.
Dynamics $f(x)$ SBF $B(x)$ Solver Method
Linear SOS SDP
PWC Dual LP, CEGIS LP, GD
Polynomial SOS SDP
PWA Inclusion SOS SDP
(Eq. (5)) PWC Dual LP, CEGIS LP, GD
StochasticBarrier.jl is a Julia toolbox for the synthesis of SBFs for System (1). The tool includes multiple independent synthesis methods with the following features:
  • Synthesis of SOS-SBFs for systems in the form of (1) with linear, polynomial and PWA inclusion dynamics,
  • Computation of transition kernel bounds for System (1) with linear and PWA inclusion dynamics,
  • Synthesis of PWC-SBFs through LP (dual and CEGIS) and GD optimization methods given the kernel bounds.
In Table 1, the SBF methods are outlined for different classes of dynamics $f$. The solver method refers to the convex optimization class. To instantiate an instance of SBF synthesis problem, the system model needs to be defined first, which consists of the dynamics and the relevant sets (safe, unsafe and initial). In Section 3.1, we show how to setup these system related parameters. Then, the SBF synthesis method needs to be specified, as described in Section 3.2.
3.1. System Setup
StochasticBarrier.jl dispatches and checks compatibility with the chosen synthesis algorithm based on the system type. We give examples about how to construct a linear, polynomial, and PWA inclusion system, and specify the safe, initial, and unsafe sets. The examples are for 2D-systems, but the user can easily set up an $n$-dimensional problem.
3.1.1. Linear Dynamics Example
.With linear dynamics, both SOS and PWC classes of SBFs can be used.
SOS-SBF
To synthesize a SOS-SBF for System (1) with linear vector field $f(x)=Ax+b$, we first need to specify matrix $A$ and bias vector $b$. The noise distribution is assumed to be zero-mean Gaussian $p_{\mathbf{w}}=\mathcal{N}(0,\sigma^{2}I)$ characterized by standard deviation $\sigma$. An example setup is shown below for a 2D-linear system.
# System Linear: SOS
A = [0.95 0.0; 0.0 0.95], b = [0.0, 0.0], $\sigma$ = [0.1, 0.1]
space = Hyperrectangle(low=[-1.0,-2.0], high=[1.0,2.0])
system = AdditiveGaussianLinearSystem(A, b, $\sigma$, space)
The Hyperrectangle struct of LazySets.jl, a Julia package for calculus with convex sets (Forets and Schilling, 2021), is used to define the state space. A struct of AdditiveGaussianLinearSystem is passed as a system object into StochasticBarrier.jl.
PWC-SBF
Using PWC-SBF requires an additional parameter to define a grid partitioning of the state space. The tool uses this grid to define the domain of PWC function in Eq. (6) and compute the transition kernel bounds in Eq. (7). The grid is parameterized by $\epsilon\in\mathbb{R}^{n}$, where the $i$-th entry is the grid cell’s half-width in the $i$-th dimension. The generate_partitions() function takes the state space and $\epsilon$ as inputs, and returns the partitions. The system object AdditiveGaussianLinearSystem, along with the state partitions, are parsed into transition_probabilities(), which is a function in StochasticBarrier.jl that computes the transition kernel bounds. An example 2D-linear system setup is provided.
# System Linear: PWC
A = [0.50 0.0; 0.0 0.50], b = [0.0, 0.0], $\sigma$ = [0.1, 0.1]
space = Hyperrectangle(low=[-1.0,-1.0], high=[0.5,0.5])
# Half-width vector
$\epsilon$ = [0.2, 0.2]
# Compute probability data
system = AdditiveGaussianLinearSystem(A, b, $\sigma$)
partitions = generate_partitions(space, $\epsilon$)
probabilities = transition_probabilities(system, partitions)
In addition, the toolbox supports saving and loading transition probability bounds to the Network Common Data Form (NetCDF) (Rew and Davis, 1990). See Appendix C for more details on this feature.
3.1.2. Polynomial Dynamics Example
.For polynomial systems, StochasticBarrier.jl only supports the SOS class of SBFs. To set up, first define a vector of symbol variables $x$, using the @polyvar macro of DynamicPolynomials.jl (Legat, 2022). Then, the polynomial function $f$ is defined symbolically. The programmatic construct of AdditiveGaussianPolySystem is passed as an object to the barrier synthesis method. An example is provided below for a 2D system.
# System Polynomial: SOS
@polyvar x[1:2] # dim = 2
$\tau$ = 0.1
f = [x[1] + $\tau$ * x[2], x[2] + $\tau$ * (-x[1] + (1 - x[1])^2)]
$\sigma$ = [0.02, 0.02]
space = Hyperrectangle(low=[-6.0, -6.0], high=[6.0, 6.0])
system = AdditiveGaussianPolySystem(f, $\sigma$, space)
3.1.3. PWA Inclusion Dynamics Example
.Setting up the system model for PWA inclusion dynamics requires PWA bounds on dynamics $f$ of System (1), as defined in Eq. (5). These bounds can be generated using Linear Bound Propagation (Zhang et al., 2018), and stored in an NetCDF file. These files can be loaded into the tool, which offers the options of using SOS- or PWC-SBF methods.
SOS-SBF
Given a NetCDF file that contains the PWA bounds, the setup for the SOS-SBF option is as follows. First, the system flag and filename need to be defined, which together specify the path to the PWA bounds. Then, the load_dynamics() function reads this data. The bounds dataset, along with the standard deviation $\sigma$, are passed into the AdditiveGaussianUncertainPWASystem struct. See the example below for a 2D pendulum model.
# System
system_flag = "pendulum"
filename = "path/to/\$system_flag/data.nc"
dataset = open_dataset(filename)
data = load_dynamics(dataset)
$\sigma$ = [0.01, 0.01]
system = AdditiveGaussianUncertainPWASystem(data, $\sigma$)
PWC-SBF
For the PWC-SBF option, the file and system specification are similar. In addition, the transition probability data must be computed using transition_probabilities().
# Compute probability data
probabilities = transition_probabilities(system)
3.1.4. Specifying Initial and Unsafe Sets
To complete the system model setup, the initial set and unsafe region(s) need to be defined. This is done by using the Hyperrectangle struct. In the toolbox, the region $\mathbb{R}^{n}\setminus X$ is considered unsafe by default. The user can further declare unsafe regions in the domain. If more than one unsafe region is defined, the UnionSet function of LazySets.jl can be used to take their union, as shown in the example below.
initial = Hyperrectangle(low=[-0.2, -0.3], high=[0.0, 0.1])
unsafe1 = Hyperrectangle(low=[-0.55, 0.4], high=[-0.45, 0.6])
unsafe2 = Hyperrectangle(low=[0.45, 0.4], high=[0.55, 0.6])
unsafe_regions = UnionSet(unsafe1, unsafe2)
If domain $X$ does not contain any unsafe regions, the variable unsafe_regions must be declared empty: EmptySet(dim), where dim is the dimension of the system. With that, the system model setup is concluded, and one can call on various SBF synthesis methods, as explained below.
3.2. SBF Synthesis
All methods for synthesizing SBFs are called under the function synthesize_barrier(alg, system, initial_region, unsafe_region; time_horizon=N), which is designed with a unified interface of input parameters. Julia’s multiple dispatch will select the appropriate method based on the alg parameter, and each algorithm returns a result object, which contains, in addition to the barrier, the resulting safety certificate and computation time. The barrier type is determined by the algorithm. Specifying the time horizon is optional, with the default value set to $N=1$.
3.2.1. SOS-SBF
.To synthesize an SOS-SBF, synthesize_barrier is called with input SumOfSquaresAlgorithm() as the first parameter. The second parameter, the system model object defined in Section 3.1, can be any of the three system types.
N = 10
# Optimize: baseline 1 (sos approach)
res = synthesize_barrier(SumOfSquaresAlgorithm(), system,
initial_region, unsafe_region; time_horizon = N)
# Result
barr = barrier(res)
$\eta$, $\beta$, ps = eta(res), beta(res), psafe(res, N)
computation_time = total_time(res)
The optimization returns a struct of values, which include the SOS-SBF $B$, $\eta$, and $\beta$ values, and the computation time. The algorithm SumOfSquaresAlgorithm() accepts keyword arguments to configure, e.g. the underlying SDP solver and the degree of the SOS barrier. Below we list the default values and how to construct the object for customization.
alg = SumOfSquaresAlgorithm(;
barrier_degree = 4
lagrange_degree = 2
sdp_solver = Mosek.Optimizer)
Setting SumOfSquaresAlgorithm(barrier_degree = 30), for example, overrides the barrier degree. The default SDP solver for SOS optimization is Mosek (ApS, 2019).
3.2.2. PWC-SBF
For PWC-SBFs, synthesize_barrier is called with DualAlgorithm() for the dual LP approach, CEGISAlgorithm() for the CEGIS approach, and GradientDescentAlgorithm() for GD approach, respectively, as the first parameter. The second parameter should be the result from transition_probabilities.
N = 10
# Optimize: method 1 (dual approach)
dual = synthesize_barrier(DualAlgorithm(), probabilities,
initial_region, obstacle_region, time_horizon = N)
# Optimize: method 2 (CEGIS approach)
cegis = synthesize_barrier(CEGISAlgorithm(), probabilities,
initial_region, obstacle_region; time_horizon = N)
# Optimize: method 3 (GD approach)
gd = synthesize_barrier(GradientDescentAlgorithm(), probabilities,
initial_region, obstacle_region; time_horizon = N)
Similar to SOS, each algorithm returns a result struct, which, in addition to the PWC-SBF $B$, values for $\eta$, $\beta$ and the computation time, also includes $\beta_{i}$, which is an upper bound on Eq. (2c) in each region $i$. Next, we highlight the core properties of each algorithm.
Linear Programming
As outlined in Section 2.2.2, amongst the LP classes, there is the dual LP method, and the CEGIS method which uses inner and outer LP optimizations. The dual LP method only has the solver as a configurable parameter.
alg = DualAlgorithm(;
linear_solver = HiGHS.Optimizer)
The default LP solver across the toolbox is HiGHS solver (Huangfu and Hall, 2018), but is customizable to any LP solver supported by JuMP, Julia Mathematical Programming (Lubin et al., 2023).
The CEGISAlgorithm in addition be configured with the number of iterations and whether
alg = CEGISAlgorithm(;
linear_solver = HiGHS.Optimizer
num_iterations = 10
distribution_guided = true)
The default is the distribution guided approach, where in each iteration one LP generates a candidate PWC-SBF that optimizes the unsafety probability given a set of finite feasible distributions. Then, the second LP generates distribution witnesses that violate the safety probability guarantee of the candidate PWC-SBF. By default, it runs for a fixed number of iterations given by num_iterations (default = 10), or it can continue adaptively until no further counterexamples exist, guaranteeing finite-time convergence (Mazouz et al., 2024a).
Projected Gradient Descent
The GD method solves the same problem as the LP approaches, utilizing the minimax formulation to define a convex loss function. The projected gradient descent method is subject to hyperparameters.
alg = GradientDescentAlgorithm(;
num_iterations = 10000
initial_lr = 1e-2
decay = 0.9999
momentum = 0.9)
The struct GradientDescentAlgorithm defines key parameters for running GD. The num_iterations parameter controls the maximum number of iterations (default: 10,000). The initial_lr sets the initial learning rate for the algorithm (default: 0.01), determining the size of each update step. The decay parameter (default: 0.9999) reduces the learning rate over time to fine-tune convergence. Finally, momentum (default: 0.9) helps to accelerate convergence by incorporating a portion of the previous gradient step into the current. Importantly, at termination, the values for $\eta$, $\beta$ are guaranteed to be sound with respect to the returned barrier, although not absolutely optimal in the sense of Theorem 2.3.
StochasticBarrier.jl is also equipped with an independent interface to easily run benchmarks. See Appendix A for details.
Table 2. Linear, polynomial and PWA inclusion stochastic dynamic systems for benchmarking in StochasticBarrier.jl.
Dynamics Name Dimension
Linear Contraction Map 1 & 2 2
Two Tank (Ramos and Dos Santos, 2007) 2
Room Temperature (Wooding et al., 2024) 3
Quadrotor (Mazouz et al., 2024a) 6
Polynomial Thermostat (Nejati et al., 2021) 1
Van der Pol Oscillator (Abate et al., 2020) 2
PWA Inclusion Pendulum NNDM (Mazouz et al., 2022) 2
Unicycle (Mazouz et al., 2024a) 4
4. Benchmark Studies
In this section, we present benchmark results for various stochastic systems with linear, polynomial and PWA inclusion dynamics. We performed two main comparison studies. First, the SOS-SBF algorithm in StochasticBarrier.jl is compared against the state-of-the-art SOS toolboxes: StochasticBarrierFunctions (Santoyo et al., 2021) (MATLAB) and PRoTECT (Wooding et al., 2024) (Python). The second evaluates all the StochasticBarrier.jl methods (SOS, DUAL-LP, CEGIS-LP, and GD) and compares their performance. All experiments are performed on a computer with a 3.9 GHz 8-Core CPU and 128 GB of memory. All methods use Mosek as the underlying SDP solver. The systems considered are summarized in Table 2 and details of the systems and their setups are provided in Appendix B. For all benchmarks, time horizon $N=10$ is used. The tool is available at https://github.com/aria-systems-group/StochasticBarrier.jl.
(a) Contraction Map
(b) Two Tank
(c) Room Temperature
(d) Quadrotor
(e) Thermostat
(f) Oscillator
Figure 1. Polynomial systems benchmarking of StochasticBarrier.jl vs StochasticBarrierFunctions (Santoyo et al., 2021) and PRoTECT (Wooding et al., 2024). Plots depict optimization time $\tau(s)$ and probability of safety $P_{s}$. If data points are not plotted beyond a certain barrier degree, it means the algorithm ran out of memory.
Comparison Against Other SOS-SBF Tools. Fig. 1 shows StochasticBarrier.jl’s performance of the SOS-SBF algorithm against StochasticBarrierFunc. (Santoyo et al., 2021) and PRoTECT (Wooding et al., 2024). A detailed table of results can be found in Appendix D. Note that as StochasticBarrierFunc. only supports linear systems, and PRo-
TECT
only linear and polynomial systems, we could only compare the tools with systems of this dynamics class. We also note that the results from PRoTECT are random, and we present average results over 10 runs. We also note that $P_{s}$ defined in Eq. (3) is computed differently in PRoTECT; we adjust it here to match our setting. This paper solely focuses on optimization-based SOS methods, i.e., disabling the objective is fundamentally different and not considered.
Across all benchmarks, StochasticBarrier.jl outperforms both StochasticBarrierFunc. and PRoTECT in terms of computation time and scalability, and tightness of safety probability. Stocha-
sticBarrierFunc.
fails to produce non-trivial results and cannot handle higher-degree SOS-SBFs (beyond degree 12), often yielding a trivial $P_{s}\geq 0$ or exceeding available memory. PRoTECT performs better but remains significantly slower (by factors ranging from $5\times$ to $1000\times$) and often returns trivial safety bounds for complex systems. StochasticBarrier.jl, in contrast, scales efficiently to high-degree SOS-SBFs (up to 30), achieves non-trivial or high safety probabilities, and maintains orders-of-magnitude faster computation across all tested models (contraction map 1, two tank, room temperature, quadrotor, thermostat, and oscillator).
(a) Contraction Map
(b) Pendulum
(c) Unicycle
Figure 2. Benchmark plots for the three PWC-SBF methods. For the pendulum and unicycle, the SOS results are also depicted for comparison. Plots depict total optimization time $\tau(s)$ and probability of safety $P_{s}$.
Benchmark Benchmark
Comparisons of StochasticBarrier.jl engines. Benchmark results for the three PWC-SBF methods implemented in StochasticBarrier.jl are summarized in Fig.2, with detailed results presented in Table 4 of Appendix D. The results show that PWC-SBFs consistently outperform SOS-SBFs, particularly for the contraction map and hybrid nonlinear systems. For the linear contraction map, the GD method achieves the fastest computation, especially as the number of partitions increases, while Dual LP and CEGIS yield tighter safety probability bounds that converge with higher partition counts. For PWA inclusion systems, PWC-SBFs are up to two orders of magnitude faster, and for the unicycle model, they provide non-trivial safety bounds where SOS fails. As expected, the Dual LP and CEGIS methods deliver the tightest bounds at the cost of higher computation time.
5. Conclusion
We introduced StochasticBarrier.jl, a Julia toolbox for generating barrier certificates and estimating safety probabilities of discrete-time stochastic dynamical systems. Benchmark results show that StochasticBarrier.jl is orders of magnitude faster and provides tighter safety bounds than existing MATLAB and Python implementations, with superior scalability to higher-dimensional systems and higher-degree SOS barriers. The toolbox also supports piecewise constant barriers, which consistently outperform SOS formulations. Current limitations include its focus on additive Gaussian noise. Future extensions will address non-Gaussian and non-additive noise.
Appendix A Benchmarking Setup
StochasticBarrier.jl is also equipped with an independent interface to easily run benchmarks. Fig. 3 depicts the general structure of the interface. The input is a .yaml file, in which the dynamics, the sets (safe, unsafe and initial), and the SBF synthesis method are specified. The output includes the SBF $B$, values for $\eta$ and $\beta$, the lower bound on safety probability $P_{s}$, and the optimization time.
Figure 3. Input-output overview of StochasticBarrier.jl.
Input and output overview of StochasticBarrier.jl interface. Graphic overview of input \$vs\$ output of StochasticBarrier.jl.
Below, we show the .yaml input files for the systems discussed above for SOS-SBF and PWC-SBF.
system_flag: "linear"
dim: 2
A: [[0.95, 0.0], [0.0, 0.95]]
b: [0.0, 0.0]
$\sigma$: [0.01, 0.01]
state_space:
low: [-1.0, -2.0]
high: [1.0, 2.0]
barrier_settings:
barrier_type: "SOS"
time_horizon: 10
barrier_degree: 12
system_flag: "polynomial"
dim: 2
f: ["0.1*x[2]+x[1]", "0.1+x[2]-0.3*x[1]+0.1*x[1]^2"]
$\sigma$: [0.02, 0.02]
state_space:
low: [-6.0, -6.0]
high: [6.0, 6.0]
initial_region:
low: [-5.0, -5.0]
high: [5.0, 5.0]
barrier_settings:
barrier_type: "SOS"
time_horizon: 10
barrier_degree: 6
system_flag: "nonlinear"
dynamics: "bounds/pendulum/dynamics_120.nc"
probabilities: "data_480_sigma_[0.01, 0.01].nc"
dim: 2
$\sigma$: : [0.01, 0.01]
initial_region:
c: [0.0, 0.0]
r: [0.01]
barrier_settings:
barrier_type: "PWC"
optimization_type: "DUAL_ALG"
time_horizon: 10
The input parameters can be changed according to the scope of the problem. For PWC-SBF, the options for the barrier optimization_type are: ”DUAL_ALG”, ”CEGIS_ALG”, or ”GD_ALG”.
This interface is primarily intended to serve as a means for utilizing the toolbox’s capabilities. However, the modular design ensures that StochasticBarrier.jl can be seamlessly integrated into larger frameworks or other toolboxes. Such integration enables researchers and developers to leverage its features within broader systems.
Appendix B Benchmark Systems
B.1. Linear Systems
Contraction Map
We consider two settings of contractive linear stochastic system in $\mathbb{R}^{2}$. For the first setting, used for the SOS benchmarks, the dynamics (Contraction Map 1) are
$\mathbf{x}(k+1)=0.95I\ \mathbf{x}(k)+\mathbf{w},$
where $I$ is the identity matrix and $\mathbf{w}\sim\mathcal{N}(0,10^{-2}I)$. The domain and initial sets are $X=[-1,2]^{2}$ and $X_{0}=[-0.05,0.05]^{2}$, respectively.
For the second set of benchmarks (PWC vs SoS engines), the Contraction Map 2 dynamics are
$\mathbf{x}(k+1)=0.50I\ \mathbf{x}(k)+\mathbf{w},$
using the same distribution $\mathbf{w}\sim\mathcal{N}(0,10^{-2}I)$. The domain and initial sets are $X=[-1,0.5]^{2}$ and $X_{0}=[-0.05,0.05]^{2}$, respectively.
Two Tank
Next, we consider a stochastic discrete-time two tank model from (Ramos and Dos Santos, 2007), with dynamics
$\mathbf{x}_{1}(k+1)$ $=\left(1-\tau\frac{\alpha_{1}}{A_{1}}\right)\mathbf{x}_{1}(k)+\tau\frac{q_{1}}{A_{1}}+\mathbf{w}_{1},$
$\mathbf{x}_{2}(k+1)$ $=\tau\frac{\alpha_{1}}{A_{2}}\mathbf{x}_{1}(k)+\left(1-\tau\frac{\alpha_{2}}{A_{2}}\right)\mathbf{x}_{2}+\tau\frac{q_{0}}{A_{2}}+\mathbf{w}_{2}(k),$
where $\mathbf{x}$ is the fluid heights in the two tanks and $\mathbf{w}\sim\mathcal{N}(0,10^{-4}I)$ the additive noise. The system parameters are $\tau=0.1$, $\frac{\alpha_{1}}{A_{1}}=1$, $\frac{q_{1}}{A_{1}}=4.5$, $\frac{\alpha_{1}}{A_{2}}=1$, $\frac{\alpha_{2}}{A_{2}}=1$, and $\frac{q_{o}}{A_{2}}=-3$. The domain and initial sets are $X=[1,9]^{2}$ and $X_{0}=[2.75,3.25]^{2}$, respectively.
Room Temperature model
The 3D room temperature model described in (Wooding et al., 2024) has deterministic dynamics
$\mathbf{x}_{1}(k+1)$ $=(1-\tau(\alpha+\alpha_{e}))\mathbf{x}_{1}(k)+\tau\alpha\mathbf{x}_{2}(k)+\tau\alpha_{e}Te$
$\mathbf{x}_{2}(k+1)$ $=(1-\tau(2\alpha+\alpha_{e}))\mathbf{x}_{2}(k)+\tau\alpha(\mathbf{x}_{1}(k)+\mathbf{x}_{3}(k))$
$\hskip 14.22636pt+\tau\alpha_{e}Te$
$\mathbf{x}_{3}(k+1)$ $=(1-\tau(\alpha+\alpha_{e}))\mathbf{x}_{3}(k)+\tau\alpha\mathbf{x}_{2}(k)+\tau\alpha_{e}Te$
to which noise $\mathbf{w}\sim\mathcal{N}(0,10^{-2}I)$ is added. The states $\mathbf{x}_{1},\mathbf{x}_{2}$, and $\mathbf{x}_{3}$ define the temperature of each room. The ambient temperature $T_{e}=10^{\circ}$C, while the heat coefficients are $\alpha_{e}=8\cdot 10^{-3}$ and $\alpha=6.2\cdot 10^{-3}$. The sampling time $\tau$ = 5 minutes. Regions of interest are $X=[17,29]^{3}\text{ and }X_{0}=[18,19]^{3}$.
Quadrotor
We consider the lateral quadrotor model from (Mazouz et al., 2024a)
$\Delta\dot{y}_{E}=\Delta v,$ $\Delta\dot{v}=g\Delta\phi,$
$\Delta\dot{\phi}=\Delta p,$ $\Delta\dot{p}=I_{x}^{-1}\Delta L_{c},$
where $y_{E}\in[-0.5,2.0]$ is the $y$-position, and $v\in[-1.0,1.0]$ is the corresponding velocity, $g$ is the gravitational acceleration, $I_{x}$ is the inertia about the $x$-axis, $\phi\in[-0.1,0.1]$ is the roll angle, $p\in[-0.1,0.1]$ is the roll rate, and $L_{c}$ is the roll control moment. The lateral model is appended with guidance for vertical motion
$\Delta\dot{z}_{E}=\Delta w,\qquad\Delta\dot{w}=\frac{1}{m}\Delta Z_{c},$
where $m$ denotes the quadrotor mass, $z_{E}\in[-0.5,3.0]$ denotes the $z$-position and $w\in[-0.5,1.5]$ is the yaw rate. The initial set is
$\ell$ $=[-0.01,0.01,-0.01,0.01,-0.01,0.01],$
$u$ $=[0.01,0.03,0.01,0.03,0.01,0.03],$
$X_{0}$ $=\{\,x\in\mathbb{R}^{6}\;\mid\;\ell\leq x\leq u\,\}.$
The control law follows directly from a stabilizing LQR controller at the nominal point. Using Euler’s method with $\Delta t=0.01$, a discrete-time model is obtained, to which $\mathbf{w}\sim\mathcal{N}(0,10^{-6}I)$ is added.
B.2. Polynomial Systems
Thermostat Model
A polynomial room temperature module taken from (Nejati et al., 2021). The stochastic difference equation is given by
$\mathbf{x}_{k+1}=(1-\beta-\theta\nu)\mathbf{x}_{k}+\theta T_{h}\nu+\beta T_{e}+\mathbf{w},$
where $\mathbf{x}_{k}$ is the room temperature, the ambient temperature $T_{e}=-15^{\circ}$C, and the heater temperature $T_{h}=45^{\circ}$C. Conduction factors are given as $\theta=0.145,\beta=0.6$, and $\nu=-0.0120155\mathbf{x}_{k}+0.8$. The system evolves over the state space $X=[1,50]$ with the initial set $X_{0}=[19.5,20]$, and unsafe regions are $X_{u}=[1,17]\cup[23,50]$. The noise parameter is $\mathbf{w}\sim\mathcal{N}(0,\cdot 10^{-2}I)$.
Van der Pol Oscillator
The Van Der Pol oscillator is from (Abate et al., 2020) has the following dynamics
$\mathbf{x}_{1}(k+1)$ $=\mathbf{x}_{1}(k)+\tau\mathbf{x}_{2}(k)+\mathbf{w}_{1},$
$\mathbf{x}_{2}(k+1)$ $=\mathbf{x}_{2}(k)+\tau(-\mathbf{x}_{1}(k)+(1-\mathbf{x}_{1}(k)^{2})\mathbf{x}_{2}(k))+\mathbf{w}_{2}.$
In this model, $\tau=0.1$ is the sampling time. The domain of operation is defined by $X=[-6,6]^{2}$ and the initial set by $X_{0}=[-5,5]^{2}$. The noise parameter is $\mathbf{w}\sim\mathcal{N}(0,4\cdot 10^{-4}I)$.
B.3. PWA Inclusion Systems
Pendulum NNDM
In this case study, we consider the pendulum NNDM model from (Mazouz et al., 2022). The dynamics are $\mathbf{x}=f^{NN}(x)+\mathbf{w}$, where $f^{NN}$ is trained on a Pendulum agent in a closed-loop with a NN controller. For this system, noise term $\mathbf{w}\sim\mathcal{N}(0,10^{-4}I)$. The control limit are $u\in[-1,1]$ and the dynamics of the pendulum are
$\dot{\theta}_{k+1}$ $=\dot{\theta}_{k}+\frac{3g}{2l}\sin(\theta_{k})\delta t^{2}+\frac{3}{ml^{2}}u\delta t^{2},$
$\theta_{k+1}$ $=\theta_{k}+\dot{\theta}_{k+1}\delta t.$
The initial set for this sytem is $X_{0}=[0.0,0.01]^{2}.$
Hybrid Unicycle
Finally, we considered the wheeled mobile robot example from (Mazouz et al., 2024a) with the dynamics of a unicycle:
$\dot{x}=v\cos\theta,\quad\dot{y}=v\sin\theta,\quad\dot{\theta}=\omega,\quad\dot{v}=a,$
where $x\in[-1.0,0.5]$ and $y\in[-0.5,1.0]$ are the Cartesian positions. The angle $\theta\in[-1.75,0.5]$ is the orientation with respect to the $x$-axis, and $v\in[-0.5,1.0]$ is the speed of the vehicle. The initial set is defined by a hyperrectangle, where $x\in[-0.01,0.01]^{4}$. The inputs are steering rate $\omega$ and acceleration $a$. The combined feedback linearization and LQR stabilizer make this system hybrid. The Euler method with $\Delta t=0.01$ defines the discrete-time dynamics, to which noise term $\mathbf{w}\sim\mathcal{N}(0,10^{-4}I)$ is added.
Appendix C Saving and loading of probability data
Our toolbox supports storing transition probability bounds and PWA bounds for non-linear functions as NetCDF, which allows for efficient storage and retrieving of multidimensional data in addition to axis annotation and attributes (Rew and Davis, 1990). First, to save the transition data, one can run the following to first create a NetCDF dataset and saving it.
# Save to a .nc file
filename = "probability_data.nc"
dataset = create_probability_dataset(probabilities) # NetCDF dataset
savedataset(dataset, path=filename;
driver=:netcdf, overwrite=true, compress=1)
The process in reverse is to first load the NetCDF dataset followed by extracting and transforming data to the appropriate format via the load_probabilities function.
filename = "set/path/to/file.nc"
dataset = open_dataset(filename)
probabilities = load_probabilities(dataset)
To load PWA bounds, use the load_dynamics(dataset) function, which expects in the .nc file a num_regions property as well as 3 datasets: nominal_dynamics_A, nominal_dynamics_b, and regions. The dataset regions should have three axes: region, dir, and x of dimension 1 to num_regions, ["lower", "upper"], and 1 to dim, respectively. nominal_dynamics_A and nominal_dynamics_b are similar with axes region, dir, y, x and region, dir, y, respectively. Ordering of axes is irrelevant; the tool will select the axis based on their name. Furthermore, the tool executes a number of consistency checks and will throw an informative error about the inconsistency if any.
Appendix D Table Results
Detailed results for the SOS and PWC benchmarks are presented in Table 3 and 4, respectively.
Table 3. Benchmark SOS methods: StochasticBarrier.jl vs MATLAB and Python implementations for linear and polynomial systems. All verifications are performed for $\mathbf{N=10}$ time-steps. OM denotes out-of-memory, while FAILED indicates that the algorithm terminated without a valid barrier. The shaded areas indicate that the toolbox cannot accommodate these systems.
Class System Deg. StochasticBarrierFunc. (Santoyo et al., 2021) PRoTECT (Wooding et al., 2024) StochasticBarrier.jl
$\tau$ (s) $\eta$ $\beta$ $P_{s}$ $\tau$ (s) $\eta$ $\beta$ $P_{s}$ $\tau$ (s) $\eta$ $\beta$ $P_{s}$
Linear 2D Contraction Map 1 2 23.56 1.00 0.00 0.00 0.58 $1.7e^{-2}$ $1.8e^{-2}$ 0.80 0.01 $1.7e^{-2}$ $1.8e^{-2}$ 0.80
4 74.86 1.00 0.00 0.00 1.51 $3.7e^{-3}$ $5.7e^{-3}$ 0.94 0.02 $3.7e^{-3}$ $5.7e^{-3}$ 0.94
8 543.29 1.00 0.00 0.00 12.85 $1.1e^{-3}$ $1.8e^{-3}$ 0.98 0.05 $1.1e^{-3}$ $1.8e^{-3}$ 0.98
12 2582.17 1.00 0.00 0.00 76.88 $1.0e^{-3}$ $1.3e^{-3}$ 0.99 0.21 $1.1e^{-3}$ $1.3e^{-3}$ 0.99
24 OM $\times$ $\times$ $\times$ 2442.43 $8.3e^{-6}$ $9.5e^{-8}$ 0.99 18.08 $9.9e^{-4}$ $1.2e^{-3}$ 0.99
30 OM $\times$ $\times$ $\times$ 9569.64 $4.6e^{-7}$ $7.8e^{-9}$ 0.99 107.14 $9.6e^{-4}$ $1.2e^{-3}$ 0.99
2D Two Tank 4 90.87 1.00 0.00 0.00 2.37 $1.00$ $3.6e^{-6}$ 0.00 0.14 $8.7e^{-2}$ $4.1e^{-2}$ 0.50
6 283.36 1.00 0.00 0.00 9.48 FAILED FAILED 0.00 0.10 $2.9e^{-2}$ $7.9e^{-5}$ 0.97
8 884.79 1.00 0.00 0.00 53.60 FAILED FAILED 0.00 0.16 $6.8e^{-4}$ $4.2e^{-5}$ 1.00
3D Room Temp. 2 109.74 FAILED FAILED 0.00 1.38 FAILED FAILED 0.00 0.12 FAILED FAILED 0.00
4 OM $\times$ $\times$ $\times$ 39.02 FAILED FAILED 0.00 0.27 $9.7e^{-2}$ $7.6e^{-3}$ 0.83
6 OM $\times$ $\times$ $\times$ 2422.07 FAILED FAILED 0.00 11.08 $9.0e^{-7}$ $8.6e^{-8}$ 1.00
8 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 20.02 $8.7e^{-6}$ $2.1e^{-6}$ 1.00
12 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 3716.59 $4.1e^{-8}$ $9.7e^{-9}$ 1.00
6D Quadrotor 2 OM $\times$ $\times$ $\times$ 9587.45 FAILED FAILED 0.00 0.14 $1.0e^{-1}$ $5.9e^{-4}$ 0.89
4 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 3.35 $9.0e^{-4}$ $1.0e^{-5}$ 1.00
6 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 148.65 $8.8e^{-6}$ $1.7e^{-7}$ 1.00
Polynomial 1D Thermostat 4 0.59 FAILED FAILED 0.00 0.05 0.87 $3.3e^{-3}$ 0.09
6 0.79 FAILED FAILED 0.00 0.08 $5e^{-11} 3.3$$3e^{-12}$ 1.00
8 1.13 FAILED FAILED 0.00 0.16 $4.2e^{-5}$ $8.5e^{-6}$ 1.00
2D Oscillator 6 21.21 FAILED FAILED 0.00 0.96 1.00 $1.8e^{-8}$ 0.00
8 168.79 FAILED FAILED 0.00 3.68 1.00 $2.9e^{-8}$ 0.00
12 5502.47 FAILED FAILED 0.00 77.03 $1.2e^{-7}$ $1.2e^{-8}$ 1.00
Table 4. Benchmark: SOS vs PWC implementations for linear and PWA inclusion systems using StochasticBarrier.jl.
All verifications are performed for $\mathbf{N=10}$ time-steps. OM denotes out-of-memory.
Class System SOS PWC DUAL LP PWC CEGIS PWC GD
$|\mathbf{Q}|$ Deg. $\tau$ (s) $P_{s}$ $|\mathbf{Q}|$ $\tau$ (s) $P_{s}$ $\tau$ (s) $P_{s}$ $\tau$ (s) $P_{s}$
Linear 2D Contraction Map 2 NA 4 4.36 0.23 64 0.44 0.99 2.43 0.99 1.57 0.95
NA 8 13.33 0.43 225 105.26 0.99 19.73 0.99 10.08 0.97
NA 20 314.22 0.88 289 248.53 0.99 39.40 0.99 34.53 0.97
NA 30 2162.92 0.90 361 2056.58 0.99 50.14 0.99 49.18 0.98
PWA Inclusion 2D Pendulum 120 2 29.10 0.93 120 0.41 0.88 1.94 0.88 2.02 0.88
240 2 43.42 0.93 240 101.38 0.94 13.70 0.94 4.81 0.94
480 2 118.45 0.94 480 208.65 0.97 44.70 0.97 18.79 0.96
4D Unicycle 1250 4 1819.45 0.00 1250 1608.43 0.65 122.30 0.64 14.76 0.64
1800 4 2455.66 0.00 1800 2143.80 0.65 211.51 0.65 20.07 0.64

### v2

@@
 Compared to its closest competitor, StochasticBarrier.jl is up to $1000\times$ faster for successful synthesis, improves safety probability bounds from 0 to $1.0$ in some cases, and handles systems with twice the dimensionality.
+1.1. Related Work
+Barrier certificates are synthesized to proof safety for dynamical systems for both deterministic (Schneeberger et al., 2023; Wang et al., 2018), and stochastic systems (Prajna et al., 2002; Jagtap et al., 2020).
+The latter, referred to as stochastic barrier functions (SBFs), are utilized to obtain a formal lower bound on the probability of remaining safe, starting from an initial set, under a well-defined probability measure.
+SBFs have been used to verify safety for linear and polynomial (Santoyo et al., 2021), non-linear (Mazouz et al., 2022), and data-driven systems (Mazouz et al., 2024b).
+Sum-of-Squares (SOS) optimization is a popular method for SBF synthesis (Prajna et al., 2002, 2007; Santoyo et al., 2021; Jagtap et al., 2020; Mazouz et al., 2022).
+Some works utilize SOSTOOLS (Prajna et al., 2002) to obtain these SBFs.
+The main downside of SOSTOOLS-based SBFs is that the algorithms are embedded in MATLAB, which scales poorly when bridging to SDP.
+PRoTect (Wooding et al., 2024), a Python based implementation, scales better, but is limited to low-degree SOS barriers and systems of polynomial class.
+Our toolbox, which is implemented in Julia, provides significant speedups for intensive numerical optimization tasks and memory management (Perkel and others, 2019; Coleman et al., 2021).
+Related tools for the verification of stochastic systems in Julia include IntervalMDP.jl (Mathiesen et al., 2024a), which performs CUDA-accelerated value iteration for Interval Markov Decision Processes (IMDPs).
+JuliaReach (Bogomolov et al., 2019) performs set-based reachability for deterministic systems.
+A tool for Satisfiability modulo theories, Satisfiability.jl (Soroka et al., 2023), provides high-level representations for SMT formulae.
+Finally, ModelVerification.jl (Wei et al., 2024) verifies DNNs and safety specifications.
+None of these verify safety for stochastic systems using SBFs.
+A platform for this purpose is provided in StochasticBarrier.jl.
+The remainder of the paper is organized as follows.
+In Section 3, we provide a detailed overview of the toolbox, with instructions and examples on setting up system models and invoking SBF synthesis engines.
+In Section 4, we demonstrate the performance of StochasticBarrier.jl through a series of case studies and benchmark comparisons.
+Finally, we conclude the paper with a summary and closing remarks in Section 5.
 2. Theoretical Framework
 We provide an overview of the SBF theoretical framework.
+An outline of the problem formulation is provided, followed by a review of the implemented methods.
+For more detail, we refer the reader to (Mazouz et al., 2022, 2024a).
 2.1. Problem
@@
 StochasticBarrier.jl implements all three methods for System (1) with linear and PWA inclusion dynamics.
+The toolbox internally computes the kernel bounds in Eq.
+(7).
 3. Overview of StochasticBarrier.jl
@@
 Plots depict optimization time $\tau(s)$ and probability of safety $P_{s}$.
-If data points are not plotted beyond a certain barrier degree, it means the algorithm ran out of memory.
+If data points are not plotted, it means the algorithm ran out of memory.
+t Refer to caption (a) Contraction Map Refer to caption (b) Pendulum Refer to caption (c) Unicycle
+Figure 2.
+Benchmark plots for the three PWC-SBF methods.
+For the pendulum and unicycle, the SOS results are also depicted for comparison.
+Plots depict total optimization time $\tau(s)$ and probability of safety $P_{s}$.
 Comparison Against Other SOS-SBF Tools.
@@
 StochasticBarrier.jl, in contrast, scales efficiently to high-degree SOS-SBFs (up to 30), achieves non-trivial or high safety probabilities, and maintains orders-of-magnitude faster computation across all tested models (contraction map 1, two tank, room temperature, quadrotor, thermostat, and oscillator).
-(a) Contraction Map
-(b) Pendulum
-(c) Unicycle
-Figure 2.
-Benchmark plots for the three PWC-SBF methods.
-For the pendulum and unicycle, the SOS results are also depicted for comparison.
-Plots depict total optimization time $\tau(s)$ and probability of safety $P_{s}$.
-Benchmark Benchmark
 Comparisons of StochasticBarrier.jl engines.
-Benchmark results for the three PWC-SBF methods implemented in StochasticBarrier.jl are summarized in Fig.2, with detailed results presented in Table 4 of Appendix D.
+Benchmark results for the three PWC-SBF methods implemented in StochasticBarrier.jl are summarized in Fig.
+2, with detailed results presented in Table 4 of Appendix D.
 The results show that PWC-SBFs consistently outperform SOS-SBFs, particularly for the contraction map and hybrid nonlinear systems.
@@
 The input is a .yaml file, in which the dynamics, the sets (safe, unsafe and initial), and the SBF synthesis method are specified.
+The output includes the SBF $B$, values for $\eta$ and $\beta$, the lower bound on safety probability $P_{s}$, and the optimization time.
 The output includes the SBF $B$, values for $\eta$ and $\beta$, the lower bound on safety probability $P_{s}$, and the optimization time.
@@
 $\mathbf{x}_{2}(k+1)$ $=\tau\frac{\alpha_{1}}{A_{2}}\mathbf{x}_{1}(k)+\left(1-\tau\frac{\alpha_{2}}{A_{2}}\right)\mathbf{x}_{2}+\tau\frac{q_{0}}{A_{2}}+\mathbf{w}_{2}(k),$
-where $\mathbf{x}$ is the fluid heights in the two tanks and $\mathbf{w}\sim\mathcal{N}(0,10^{-4}I)$ the additive noise.
+where $\mathbf{x}$ is the fluid heights in the two tanks and $\mathbf{w}\sim\mathcal{N}(0,10^{-2}I)$ the additive noise.
 The system parameters are $\tau=0.1$, $\frac{\alpha_{1}}{A_{1}}=1$, $\frac{q_{1}}{A_{1}}=4.5$, $\frac{\alpha_{1}}{A_{2}}=1$, $\frac{\alpha_{2}}{A_{2}}=1$, and $\frac{q_{o}}{A_{2}}=-3$.
-The domain and initial sets are $X=[1,9]^{2}$ and $X_{0}=[2.75,3.25]^{2}$, respectively.
+The domain and initial sets are $X=[1,10]^{2}$ and $X_{0}=[4.75,5.25]^{2}$, respectively.
 Room Temperature model
@@
 where $\mathbf{x}_{k}$ is the room temperature, the ambient temperature $T_{e}=-15^{\circ}$C, and the heater temperature $T_{h}=45^{\circ}$C.
-Conduction factors are given as $\theta=0.145,\beta=0.6$, and $\nu=-0.0120155\mathbf{x}_{k}+0.8$.
+Conduction factors are given as $\theta=0.145,\beta=0.06$, and $\nu=-0.0120155\mathbf{x}_{k}+0.8$.
 The system evolves over the state space $X=[1,50]$ with the initial set $X_{0}=[19.5,20]$, and unsafe regions are $X_{u}=[1,17]\cup[23,50]$.
@@
 $\tau$ (s) $\eta$ $\beta$ $P_{s}$ $\tau$ (s) $\eta$ $\beta$ $P_{s}$ $\tau$ (s) $\eta$ $\beta$ $P_{s}$
-Linear 2D Contraction Map 1 2 23.56 1.00 0.00 0.00 0.58 $1.7e^{-2}$ $1.8e^{-2}$ 0.80 0.01 $1.7e^{-2}$ $1.8e^{-2}$ 0.80
-4 74.86 1.00 0.00 0.00 1.51 $3.7e^{-3}$ $5.7e^{-3}$ 0.94 0.02 $3.7e^{-3}$ $5.7e^{-3}$ 0.94
-8 543.29 1.00 0.00 0.00 12.85 $1.1e^{-3}$ $1.8e^{-3}$ 0.98 0.05 $1.1e^{-3}$ $1.8e^{-3}$ 0.98
-12 2582.17 1.00 0.00 0.00 76.88 $1.0e^{-3}$ $1.3e^{-3}$ 0.99 0.21 $1.1e^{-3}$ $1.3e^{-3}$ 0.99
-24 OM $\times$ $\times$ $\times$ 2442.43 $8.3e^{-6}$ $9.5e^{-8}$ 0.99 18.08 $9.9e^{-4}$ $1.2e^{-3}$ 0.99
-30 OM $\times$ $\times$ $\times$ 9569.64 $4.6e^{-7}$ $7.8e^{-9}$ 0.99 107.14 $9.6e^{-4}$ $1.2e^{-3}$ 0.99
-2D Two Tank 4 90.87 1.00 0.00 0.00 2.37 $1.00$ $3.6e^{-6}$ 0.00 0.14 $8.7e^{-2}$ $4.1e^{-2}$ 0.50
-6 283.36 1.00 0.00 0.00 9.48 FAILED FAILED 0.00 0.10 $2.9e^{-2}$ $7.9e^{-5}$ 0.97
-8 884.79 1.00 0.00 0.00 53.60 FAILED FAILED 0.00 0.16 $6.8e^{-4}$ $4.2e^{-5}$ 1.00
-3D Room Temp. 2 109.74 FAILED FAILED 0.00 1.38 FAILED FAILED 0.00 0.12 FAILED FAILED 0.00
-4 OM $\times$ $\times$ $\times$ 39.02 FAILED FAILED 0.00 0.27 $9.7e^{-2}$ $7.6e^{-3}$ 0.83
-6 OM $\times$ $\times$ $\times$ 2422.07 FAILED FAILED 0.00 11.08 $9.0e^{-7}$ $8.6e^{-8}$ 1.00
-8 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 20.02 $8.7e^{-6}$ $2.1e^{-6}$ 1.00
-12 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 3716.59 $4.1e^{-8}$ $9.7e^{-9}$ 1.00
-6D Quadrotor 2 OM $\times$ $\times$ $\times$ 9587.45 FAILED FAILED 0.00 0.14 $1.0e^{-1}$ $5.9e^{-4}$ 0.89
-4 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 3.35 $9.0e^{-4}$ $1.0e^{-5}$ 1.00
-6 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 148.65 $8.8e^{-6}$ $1.7e^{-7}$ 1.00
-Polynomial 1D Thermostat 4 0.59 FAILED FAILED 0.00 0.05 0.87 $3.3e^{-3}$ 0.09
-6 0.79 FAILED FAILED 0.00 0.08 $5e^{-11} 3.3$$3e^{-12}$ 1.00
-8 1.13 FAILED FAILED 0.00 0.16 $4.2e^{-5}$ $8.5e^{-6}$ 1.00
-2D Oscillator 6 21.21 FAILED FAILED 0.00 0.96 1.00 $1.8e^{-8}$ 0.00
-8 168.79 FAILED FAILED 0.00 3.68 1.00 $2.9e^{-8}$ 0.00
-12 5502.47 FAILED FAILED 0.00 77.03 $1.2e^{-7}$ $1.2e^{-8}$ 1.00
+Linear 2D Contraction Map 1 2 16.48 1.00 0.00 0.00 0.69 $1.7e^{-2}$ $1.8e^{-2}$ 0.80 0.01 $1.7e^{-2}$ $1.8e^{-2}$ 0.80
+4 59.95 1.00 0.00 0.00 1.52 $3.7e^{-3}$ $5.7e^{-3}$ 0.94 0.02 $3.7e^{-3}$ $5.7e^{-3}$ 0.94
+8 487.98 1.00 0.00 0.00 12.98 $1.1e^{-3}$ $1.8e^{-3}$ 0.98 0.05 $1.1e^{-3}$ $1.8e^{-3}$ 0.98
+12 2157.99 1.00 0.00 0.00 76.57 $1.0e^{-3}$ $1.3e^{-3}$ 0.99 0.23 $1.1e^{-3}$ $1.3e^{-3}$ 0.99
+24 OM $\times$ $\times$ $\times$ 2998.25 $8.3e^{-6}$ $9.5e^{-8}$ 0.99 18.74 $9.9e^{-4}$ $1.2e^{-3}$ 0.99
+30 OM $\times$ $\times$ $\times$ 9887.87 $4.6e^{-7}$ $7.8e^{-9}$ 0.99 107.47 $9.6e^{-4}$ $1.2e^{-3}$ 0.99
+2D Two Tank 10 OM $\times$ $\times$ 289.69 FAILED FAILED 0.00 0.78 $5e^{-7} 2.2$$2e^{-8}$ 1.00
+12 OM $\times$ $\times$ $\times$ 1789.52 FAILED FAILED 0.00 1.69 $2.7e^{-9}$ $7.4e^{-10}$ 1.00
+14 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 4.49 $2.3e^{-7}$ $6.8e^{-8}$ 1.00
+3D Room Temp. 6 OM $\times$ $\times$ $\times$ 1008.67 FAILED FAILED 0.00 10.94 $3.1e^{-7}$ $2.9e^{-8}$ 1.00
+8 OM $\times$ $\times$ $\times$ 12897.12 FAILED FAILED 0.00 15.81 $3.1e^{-6}$ $7.5e^{-7}$ 1.00
+14 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 9198.78 $1.3e^{-7}$ $3.1e^{-8}$ 1.00
+6D Quadrotor 2 OM $\times$ $\times$ $\times$ 9587.45 FAILED FAILED 0.00 0.18 $1.0e^{-1}$ $5.9e^{-4}$ 0.89
+4 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 2.57 $9.0e^{-4}$ $1.0e^{-5}$ 1.00
+6 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 163.51 $8.8e^{-6}$ $1.7e^{-7}$ 1.00
+Polynomial 1D Thermostat 6 0.49 $6.9e^{-9}$ $1.0e^{-9}$ 1.00 0.10 $8.6e^{-9}$ $1.6e^{-5}$ 1.00
+8 0.89 $6.9e^{-9}$ $1.9e^{-9}$ 1.00 0.11 $1.5e^{-8}$ $2.8e^{-7}$ 1.00
+10 1.13 FAILED FAILED 0.00 0.22 $1.4e^{-8}$ $2.8e^{-7}$ 1.00
+2D Oscillator 8 168.79 FAILED FAILED 0.00 3.84 1.00 $2.9e^{-8}$ 0.00
+12 1937.42 $5.5e^{-5}$ $8.8e^{-6}$ 1.00 78.89 $1.1e^{-7}$ $1.2e^{-8}$ 1.00
+14 4872.54 FAILED FAILED 0.00 144.37 $1.2e^{-7}$ $1.2e^{-8}$ 1.00
 Table 4. Benchmark: SOS vs PWC implementations for linear and PWA inclusion systems using StochasticBarrier.jl.